OpenStudy (graciebugg):

Will Fan and Medal. Thank you in advance, Which of the following measurements could be the side lengths of a right triangle? A.48 in, 64 in, 96 in B.40 in, 64 in, 80 in C. 48 in, 72 in, 80 in D. 48 in, 64 in, 80 in

OpenStudy (anonymous):

A^2+B^2=C^2

OpenStudy (anonymous):

If it is a right triangle you should be able to use (is it called pascals rule?) the above equation

OpenStudy (anonymous):

So take the 2 smaller sides and set them equal to A or B (it doesnt matter) and solve for C. If it equals the larger side then you win!

OpenStudy (anonymous):

I dunno if you need more explanation, but i believe you can solve it :p I belieeeeeveeee

OpenStudy (graciebugg):

Ohhhhh Thank you. That makes sense :p

OpenStudy (anonymous):

We can verify answers once you do solve it tho :3

OpenStudy (graciebugg):

None of them work.

OpenStudy (graciebugg):

@q12157

OpenStudy (anonymous):

Did you try D

OpenStudy (anonymous):

48^2+64^2=?

OpenStudy (anonymous):

80^2=?

OpenStudy (graciebugg):

btw that was the Pythagorean Theorem. and I didn't square them.... Wow lol

OpenStudy (anonymous):

Oh haha well now you know. and now I know too hahah...haven't done this sort of math in yeaaarssss

OpenStudy (graciebugg):

XD I used to know this stuff. I was in third grade doing my brothers 10th grade hw.